If a new law decreases the highest marginal tax rate for all companies from 35% to 21% for next fiscal year. What would be the forecasted WACC? Market Risk premium of 4.4%. Please show formula used and work.
Answer #1

The WACC can be computed as :

WACC = weight of debt * after tax cost of debt + weight of equity * cost of equity

Cost of equity can be calculated as :

Re = Rf + beta (Rm - Rf)

As the marginal tax rate decreases, the cost of debt of companies increases which increases the WACC. So, the forecasted WACC will rise.

Earlier the after tax cost of debt was: Rd * (1 - 0.35)

= Rd * 0.65

Now, the after tax cost of debt is :

= Rd * (1 - 0.21)

= Rd * 0.79

So, the forecasted WACC rises.